# Inverse Transition Learning: Learning Dynamics from Demonstrations

### v1

@@ Appendix B Non-Convex Feasible Region
Appendix C Inference Details
HMC for Dirichlet
The log-likelihood function of the target distribution takes the following form:
$\log P(T^{*}|D)=\sum_{s,a}Dir(\mathbf{N}_{s,a}+\delta|s,a)$ (11)
Applying HMC to Dirichlet distributions is challenging due to the constrained nature of these distributions, which can hinder the efficiency of standard HMC methods. To address these issues, (Betancourt 2012) introduce a set of invertible transformations that reshape the Dirichlet distribution into a more tractable form by removing the simplex constraints of the original Dirichlet. To make our sampling more efficient in high dimensional settings, we add a logit transformation to remove the need of the Dirichlet variables to be in between 0 and 1. We will refer as $w$ as the variable obtained after applying all of these transformations on a transition dynamics $T$. Since all of the transformations are invertible we can easily go from one variable to the other. Algorithm 2 summarize how to sample from $P_{\epsilon}(T^{*}|\mathcal{D})$. We based our algorithm based on the first order leapfrog with constraint HMC algorithm of (Betancourt 2011).
HMC for Dirichlet with Reflection
Notation
  • $m$: The momentum variable, representing the current momentum.
  • $w$: The position variable, denoting the current position.
  • $T$: Corresponding transition dynamics of the position variable $w$
  • $\alpha$: The step size parameter, which determines the scale of each discrete step in the leapfrog integration.
  • $\nabla E(w)$: The gradient of the energy function at position $w$, used to compute the force in the Hamiltonian dynamics.
  • $C(T)$: All the constraints function that ensures the trajectory remains within the valid region defined by our constraints in terms of the dynamics $T$.
  • $L$: The total number of integration steps to be performed in the leapfrog algorithm.
  • $l$: An index variable representing the current time step within the loop of the leapfrog integration.
  • $w\_to\_T$ Transform variable w into variable T. (See (Betancourt 2012) )
  • $T\_to\_w$ Transform variable T into variable w. (See (Betancourt 2012) )
Note that if after the final leap frog step, the current transition dynamics $T$ does not satisfy $\textbf{constraints}(\pi^{*}(T))$ or the leap frog integration terminates outside the feasible region, it is immediately rejected. This ensures that each sample satisfies the $\epsilon$-ball property for each state $s\in\mathcal{D}$. Such “clean-up step” step is needed in case the leap frog integration stop when the position vector is outside of the constraints and did not have time to “roll-back” into the feasible region.
Algorithm 3 HMC with reflection for Bayesian ITL
1:  First momentum half step
2:  $m\leftarrow m-\frac{1}{2}\alpha\nabla E(w)$
3:  for $l=0$ to $L$ do
4:     Full spatial step
5:     $w\leftarrow w+\alpha m$
6:     Transform $w$ into $T$
7:     $T\leftarrow\text{w\_to\_T}(w)$
8:     Create $\widehat{\pi}_{\epsilon}(\cdot\mid\cdot;T^{*})$ with proposed $T$
9:     Squeeze maximum info
10:     $\pi\leftarrow\pi^{*}(T^{(i)})$
12:     if all constraints($\pi$) satisfied then
13:        Full momentum step
14:        $m\leftarrow m-\alpha\nabla E(w)$
15:     else
16:        Bounce
17:        constraints($\pi$) $c$ is violated
18:        $\hat{n}\leftarrow\nabla c(T)/\|\nabla c(T)\|$
19:        $m\leftarrow m-2(m\cdot\hat{n})\hat{n}$
20:     end if
21:  end for
22:  Full spatial step
23:  $w\leftarrow w+\alpha m$
24:  Last momentum half step
25:  $m\leftarrow m-\frac{1}{2}\alpha\nabla E(w)$
Appendix D Synthetic Environments
Gridworld Environment
@@ Appendix D Synthetic Environments
  • Soft-Wall Penalty: If the agent attempts to move across a tile designated as a soft-wall, it incurs a penalty of $-5$ reward points for each attempt. This mechanic discourages the agent from crossing these specific tiles.
  • Movement Penalty: For every other tile that the agent moves through, it receives a minor penalty of $-0.1$ reward points. This encourages the agent to find the shortest possible path to the goal.
  • Goal Reward: Upon successfully reaching the goal state, the agent is awarded $+10$ reward points. This substantial reward signifies the completion of the episode and serves as the primary incentive for the agent to navigate the grid efficiently.
@@
The updated reward function is defined identically to the original environment, with soft-wall tiles incurring a penalty of $-5$ reward points, other tiles a penalty of $-0.1$ points, and a reward of $+10$ points assigned upon reaching the goal state. The shift in the soft-wall location directly affects the agent’s trajectory, demonstrating the agent’s ability to adjust its policy in response to changes in environmental dynamics. This transfer task effectively evaluates the flexibility and robustness of the learned policy.
Randomworld Environment
We introduce the RandomWorld environment (inspired by (Jiang et al. 2015)), which is designed to evaluate the performance of reinforcement learning algorithms under conditions of high uncertainty and stochasticity. The RandomWorld environment is characterized by the following properties:
@@
  • Reward Function: Rewards are assigned randomly yet structured such that state 1 yields the highest expected reward, and state 15 the lowest. Specifically, the reward for state $s$, $R(s)$, is uniformly distributed within the interval $[16-s-1,16-s]$, aligning with the descending order of state desirability.
@@
  • Random Reward Generation: The reward for state $s$ in the transfer task, $R_{\text{transfer}}(s)$, is determined by a random draw from a uniform distribution over the range $[s-1,s]$, thus ensuring that higher state numbers correspond to higher expected rewards.
@@
Generating batch data $\mathcal{D}$
@@
Averaging Procedure for Experimental Results
@@
  • Gridworld: We generate 50 independent batch datasets for the Gridworld environment. The experimental results for each dataset are recorded, and the final result is obtained by averaging these outcomes.
  • RandomWorld: For the RandomWorld environment, we create 20 independent instances of the environment. Each of these RandomWorld instances is accompanied by 5 independently generated batch datasets, leading to a total of 100 unique datasets (20 worlds multiplied by 5 datasets each). The experimental outcomes across these datasets are compiled, and their average is computed to determine the overall performance in the RandomWorld environment.
@@
Appendix E Real-life ICU Environment
@@
About MIMIC-IV Dataset
@@
Data Preprocessing for Hypotension Analysis
@@
State Space Construction
@@
O2 Partial Pressure of Oxygen / Fraction Inspired Oxygen $<200$ and $\geq 100$ 1
O2 Partial Pressure of Oxygen / Fraction Inspired Oxygen $<100$ 2
@@
BP Mean Blood Pressure $<70$ mmHg 1
@@
GCS Glasgow Coma Scale (GCS) $>14$ 1
@@
Crea Creatinine $>1.9$ and $\leq 4.9$ mg/dL 1
Crea Creatinine $>4.9$ mg/dL 2
@@
  • Glasgow Coma Scale (GCS) ($15$): Bin 0 (since $10\leq 12$)
  • Creatinine ($2.5$ mg/dL): Bin 2 (since $5\geq 4.9$)
Thus, the tuple $(150,65,15,2.5)$ would be mapped to the discretized state $(1,1,0,2)$ according to our binning process. This discretization approach allows us to capture a comprehensive yet manageable representation of the patient’s clinical status, facilitating the application of our offline reinforcement learning model to infer the unknown dynamics $T^{*}$.
Action Space Definition
@@ 3 Both vasopressor therapy and IV fluid bolus
Reward Function Definition
@@
Transfer Task and Modified Reward Function
@@
Construction of the $\epsilon$-Optimal Expert Policy $\pi_{\epsilon}$
@@ Appendix F Results
Figure 7: ($40\%$ stochastic-policy states) Top row: Normalized Value vs. Coverage for Gridworld (left: Standard Task, middle: Transfer Task), Bottom row: Normalized Value vs. Coverage for Randomworlds (left: Standard Task, middle: Transfer Task). Rightmost plots: Normalized Value vs. Bayesian Regret of both Tasks (top: Gridworld, bottom: Randomworlds).
Table 4: Gridworld and Randomworlds Results (40% stochastic-policy states)
@@
Figure 8: (20% stochastic-policy states) Top row: Normalized Value vs. Coverage for Gridworld (left: Standard Task, middle: Transfer Task), Bottom row: Normalized Value vs. Coverage for Randomworlds (left: Standard Task, middle: Transfer Task). Rightmost plots: Normalized Value vs. Bayesian Regret of both Tasks (top: Gridworld, bottom: Randomworlds).
Table 5: Comparison of Gridworld and Randomworlds (20% stochastic-policy states)
@@
Figure 9: (0% stochastic-policy states) Top row: Normalized Value vs. Coverage for Gridworld (left: Standard Task, middle: Transfer Task), Bottom row: Normalized Value vs. Coverage for Randomworlds (left: Standard Task, middle: Transfer Task). Rightmost plots: Normalized Value vs. Bayesian Regret of both Tasks (top: Gridworld, bottom: Randomworlds).
Table 6: Comparison of Gridworld and Randomworlds (0% stochastic-policy states)

### v2

@@ Appendix A Non-Convex Feasible Region
Appendix B Learning Dynamics and Rewards Jointly
A key concern is that assuming a known reward function (beyond simple success/failure signals) is often unreasonable, particularly in settings like robotics. To address this, we extend our ITL algorithm to simultaneously learn the dynamics $T$ and rewards $R$.
This extension follows the exact same framework as our main method. However, to ensure the constraints remain linear with respect to both $R$ and $T$, we adopt an iterative approach. Specifically, for the value function terms $(I-\gamma T_{\pi})^{-1}R_{\pi}$, we fix the matrices $T_{\pi}$ and vectors $R_{\pi}$ to their values from the previous iteration (initialized with $T^{\text{MLE}}$ and $R_{\text{prior}}$). This allows us to proceed with the same convex optimization procedure as before.
The new loss function becomes:
$\min_{T,R}\sum_{s,a,s^{\prime}}N_{s,a,s^{\prime}}\left(T(s^{\prime}\mid s,a)-T^{\text{MLE}}(s^{\prime}\mid s,a)\right)^{2}+\|R-R_{\text{prior}}\|^{2}$ (12)
Here, $R_{\text{prior}}$ represents a sparse prior: 0 everywhere except at the goal state, which is assigned a value of $10$. Table 3 summarizes performance on the Gridworld environment averaged over $20$ different datasets, with $\epsilon=3$. We compare standard ITL (using the true rewards $R$) against ITL-No-R, which learns the reward jointly with the dynamics. Remarkably, ITL-No-R recovers a reward function that produces dynamics with performance comparable to ITL, despite not having access to the true rewards during training.
Table 3: Comparison of ITL (known rewards) and ITL-No-R (jointly learned dynamics and rewards) on the Gridworld environment ($\epsilon=3$).
Metric ITL ITL-No-R
Value 115.70 $\pm$ 2.02 115.37 $\pm$ 2.02
Best matching 0.65 $\pm$ 0.11 0.64 $\pm$ 0.12
$\epsilon$-matching 0.73 $\pm$ 0.14 0.73 $\pm$ 0.14
Total-Variation 132.46 $\pm$ 6.40 132.76 $\pm$ 6.44
Value Transfer 41.79 $\pm$ 8.85 41.15 $\pm$ 9.95
Best matching Transfer 0.54 $\pm$ 0.14 0.54 $\pm$ 0.13
$\epsilon$-matching Transfer 0.59 $\pm$ 0.09 0.59 $\pm$ 0.09
Time 4.82 $\pm$ 2.92 0.61 $\pm$ 0.60
Appendix C Synthetic Environments
C.1 Gridworld Environment
@@
  • Soft-Wall Penalty: If the agent attempts to move across a tile designated as a soft-wall, it incurs a penalty of 5-5 reward points for each attempt. This mechanic discourages the agent from crossing these specific tiles.
  • Movement Penalty: For every other tile that the agent moves through, it receives a minor penalty of 0.1-0.1 reward points. This encourages the agent to find the shortest possible path to the goal.
  • Goal Reward: Upon successfully reaching the goal state, the agent is awarded +10+10 reward points. This substantial reward signifies the completion of the episode and serves as the primary incentive for the agent to navigate the grid efficiently.
@@ C.1.1 Transfer Task
The updated reward function is defined identically to the original environment, with soft-wall tiles incurring a penalty of 5-5 reward points, other tiles a penalty of 0.1-0.1 points, and a reward of +10+10 points assigned upon reaching the goal state. The shift in the soft-wall location directly affects the agent’s trajectory, demonstrating the agent’s ability to adjust its policy in response to changes in environmental dynamics. This transfer task effectively evaluates the flexibility and robustness of the learned policy.
C.2 Randomworld Environment
We introduce the RandomWorld environment (inspired by (Jiang et al.,, 2015)), which is designed to evaluate the performance of reinforcement learning algorithms under conditions of high uncertainty and stochasticity. The RandomWorld environment is characterized by the following properties:
@@ C.2 Randomworld Environment
  • Reward Function: Rewards are assigned randomly yet structured such that state 1 yields the highest expected reward, and state 15 the lowest. Specifically, the reward for state $s$, $R(s)$, is uniformly distributed within the interval [16s1,16s][16-s-1,16-s], aligning with the descending order of state desirability.
@@ C.2.1 Transfer Task
  • Random Reward Generation: The reward for state $s$ in the transfer task, $R_{\text{transfer}}(s)$, is determined by a random draw from a uniform distribution over the range [s1,s][s-1,s], thus ensuring that higher state numbers correspond to higher expected rewards.
@@
C.3 Generating batch data $\mathcal{D}$
@@
C.4 Averaging Procedure for Experimental Results
@@
  • Gridworld: We generate 100 independent batch datasets for the Gridworld environment. The experimental results for each dataset are recorded, and the final result is obtained by averaging these outcomes.
  • RandomWorld: For the RandomWorld environment, we create 10 independent instances of the environment. Each of these RandomWorld instances is accompanied by 50 independently generated batch datasets, leading to a total of 500 unique datasets (10 worlds multiplied by 50 datasets each). The experimental outcomes across these datasets are compiled, and their average is computed to determine the overall performance in the RandomWorld environment.
@@
Appendix D Real-life ICU Environment
@@
D.1 About MIMIC-IV Dataset
@@
D.2 Data Preprocessing for Hypotension Analysis
@@
D.3 State Space Construction
@@
O2 Partial Pressure of Oxygen / Fraction Inspired Oxygen <200<200 and $\geq 100$ 1
O2 Partial Pressure of Oxygen / Fraction Inspired Oxygen <100<100 2
@@
BP Mean Blood Pressure <70<70 mmHg 1
@@
GCS Glasgow Coma Scale (GCS) >14>14 1
@@
Crea Creatinine >1.9>1.9 and $\leq 4.9$ mg/dL 1
Crea Creatinine >4.9>4.9 mg/dL 2
@@
  • Glasgow Coma Scale (GCS) ($10$): Bin 0 (since $10\leq 12$)
  • Creatinine ($5$ mg/dL): Bin 2 (since $5\geq 4.9$)
Thus, the tuple $(150,65,10,5)$ would be mapped to the discretized state $(1,1,0,2)$ according to our binning process. This discretization approach allows us to capture a comprehensive yet manageable representation of the patient’s clinical status, facilitating the application of our offline reinforcement learning model to infer the unknown dynamics $T^{*}$.
D.4 Action Space Definition
@@ 3 Both vasopressor therapy and IV fluid bolus
D.5 Reward Function Definition
@@
D.6 Transfer Task and Modified Reward Function
@@
D.7 How to Tune $\epsilon$:
[…]
Figure 6: ITL performance across $\epsilon$ values on a held-out dataset for the healthcare setting.
D.8 Construction of the $\epsilon$-Optimal Expert Policy $\pi_{\epsilon}$
@@ Appendix E Results
Figure 7: ($40\%$ stochastic-policy states, $\epsilon$ Gridworld = 3.0, $\epsilon$’s RandomWorlds = [2.37, 2.53, 2.21, 1.89, 4.27, 1.74, 2.37, 2.84, 2.37, 2.53]) Top row: Normalized Value vs. Coverage for Gridworld (left: Standard Task, middle: Transfer Task), Bottom row: Normalized Value vs. Coverage for Randomworlds (left: Standard Task, middle: Transfer Task). Rightmost plots: Normalized Value vs. Bayesian Regret of both Tasks (top: Gridworld, bottom: Randomworlds).
[…]
@@
[…]
@@
[…]